OpenStudy (anonymous):

Use a calculator to evaluate csc 3pie/14

OpenStudy (anonymous):

\[\frac{3\pi}{14} =\frac{3 \times \frac{22}{7}}{14} =\frac{3 \times 22}{7 \times14} =\frac{66}{98}=0.6734\]

OpenStudy (anonymous):

therefore cosec(0.6734)=0.738798290110701

OpenStudy (anonymous):

A. 1.604 B. 1.279 C. 47.686 D. 42.314 E. .739